OpenStudy (anonymous):

plissss help me help me Let s,r,l are roots of x^3-8=0. Find integer n such that s^n=r^n=l^n !

OpenStudy (anonymous):

why?

OpenStudy (a_clan):

Sorry for the hasty answer.

OpenStudy (anonymous):

oug ... pliss help me

OpenStudy (a_clan):

a^3 - 2^3 =0 Using the formula, a^3 - b^3 = (a-b)(a^2 +ab + b^2) We get, (a-2)(a^2 + 2a + 4) = 0 One of the roots is 2 (say s) But other two roots r and l are imaginary roots. \[r= [ -2 +\sqrt{-12}]/2\] \[l= [ -2 - \sqrt{-12}]/2\] We want 'n' such that s^n = r^n = l^n or \[([ -2 +\sqrt{-12}]/2)^{n}\] = s^n = 2^n But the imaginary number cannot be equal to 2. So, there cannot be any value of 'n' which satisfies the above equality.

OpenStudy (anonymous):

owh....thanks ^_^_^